### v1

@@ 2.3 Detailed Analysis on Modification with DeepBioisostere
We randomly selected three molecules from the ChEMBL database with moderate molecular weight and lipophilicity, as visualized in Fig. 2(a). We supposed scenario 1 introduced in Section 2.2, where log is to be modulated while a molecular weight is kept unchanged. We generated 100 unique modified molecular structures for each test molecule and property condition with the DeepBioisostere model. Again, only the fragments from the test dataset for insertion were adopted, the same as in the previous section. We exhibited their log values and molecular weights in Fig. 2(b). For every test molecule, the generated structures under different property conditions showed distributions that were completely distinct in terms of log values but very similar in molecular weights. log was increased or decreased by about 1 while most changes in molecular weight were less than 20, corresponding to the weight of one or two heavy atoms.
@@ 2.3.1 Removal Fragment Selection
We looked into the generated molecules and analyzed the removal fragment selection step. We counted how many times each substructure was selected as a removal fragment and visualized them in Fig. 2(c). For the molecule in the middle column, CHEMBL4071936, the carboxylic acid group was most frequently selected as a removal fragment to increase log. In contrast, to decrease log, it was hardly substituted, but the chlorodimethyl benzene moiety was preferred as a removal fragment. These choices of removal fragments align well with the known chemical knowledge; introducing a carboxylic acid group enhances hydrophilicity, whereas incorporating an aryl halide moiety increases hydrophobicity. The other two test molecules also showed similar differences in the removal fragment selection under different property control conditions, as exhibited in the 1st and 3rd columns in Fig. 2.
For the first test molecule in Fig. 2(a), CHEMBL1522368, we visualized top-3 ranked structures generated by DeepBioisostere in Fig. 2(d). When required to increase the log value (the orange arrow), our model substituted a nitrofuran ring with another fragment. All the selected three insertion fragments are aryl groups with halogen elements, which are apparently more hydrophobic than the nitrofuran ring. As a result, the provided molecular structures exhibited log values higher than the original ones: 4.26, 4.05, and 4.08. Meanwhile, when required to decrease the log value (the blue arrow), our model selected a 2,5-dimethyl phenyl group to replace with other aryl substituents. The alternative fragments contain a hydroxyl group or a pyridine moiety, resulting in the decreased log values of the modified structures: 2.20, 2.18, and 2.33. These results suggest that, as discussed in Section 2.1, our model selects insertion fragments by considering the given property control objectives and the removal fragment chosen earlier.
@@ 2.3.2 Insertion Fragment Selection - Effect of Chemical Environment on Property Control
Another essential factor that governs the selection of insertion fragments is the chemical environment. Here, the chemical environment refers to the substituents surrounding a chosen removal substructure to be alternated by a new insertion fragment. We demonstrate its importance in molecular optimization by two single property control tasks: increasing log by 1 and increasing QED by 0.2. Obviously, the latter task is more complicated to achieve than the former since QED is calculated based on eight individual molecular features, including log.
@@
Fig. 3 illustrates the overall experimental procedure and its results. First, we selected two molecules that contain a pyrazole group, which is a frequently considered aromatic group in bioisosteric replacement.[11] Their molecular structures are shown in Fig. 3(a), and we will refer to them as molecules A and B, respectively. The two molecules exhibit great differences in log(4.65 and -1.39) while showing moderate levels of QED values(0.58 and 0.60). Notably, the pyrazole moieties in the two molecules are placed in totally different chemical environments, particularly in terms of the number of aromatic rings, hydrogen bond donors and acceptors, and rotatable bonds. Here, we forced our model to choose only the pyrazole groups as a removal fragment to modulate a given target property, yielding 100 unique modified structures, respectively. (AA and BB in Fig. 3(a)) For comparison, we retrieved the insertion fragments selected for molecule A and adopted them into molecule B, resulting in another set of 100 unique molecules. (AB in Fig. 3(a)) In the first two cases, AA and BB, the individual chemical environments directly engaged in the selection of insertion fragments. However, in the case of AB, the insertion fragments were selected based solely on the chemical environment of molecule A, without any involvement of molecule B. By the comparison between those cases, we examine the significance of the chemical environment in molecular property control, especially during the insertion fragment selection.
@@
Fig. 3(b) and (c) show the property distributions of the modified molecules, where both cases of AA and BB almost achieved the objective values. However, in the case of AB, the observed property changes in log and QED conditioning tasks were distinct. In the former task, insertion fragments selected to increase log of molecule A also increased log of molecule B, despite the difference between the chemical environments that the pyrazole moieties were involved in. Conversely, in the QED conditioning task, the case of AB resulted in molecular structures with decreased QED values while those from AA and BB exhibited increased QED values compared to the original ones. This apparently shows that the bioisosteric replacements selected to improve the QED value of molecule A failed to improve that of molecule B. We attribute this distinction to the intrinsic difference between the two properties; log is readily calculated by summing over the estimated contributions of each chemical moiety in a molecule while QED is calculated with a non-linear function of the molecule-level structural features such as the numbers of hydrogen bond donors and acceptors. This example study suggests that the surrounding parts, not only the removal fragment, should be considered when choosing a suitable bioisosteric replacement to modulate a target molecular property dependent on the whole molecular structure in a broad sense. It is especially important in drug design, where crucial properties such as synthetic accessibility and ligand binding affinity depend on the non-linear relationships between internal substructures or structural features.
2.4 Finding Novel Bioisosteres for Uncommon Molecular Fragment
@@
Next, we examine our model in finding novel bioisosterism. Identification of novel bioisosteres becomes necessary when modification of an uncommon molecular fragment is required. In this case, a typical database mining approach is likely to be insufficient since only a few bioisosteres from the database might be available for that fragment. Indeed, among 206,008 fragments in the fragment library we created, 102,467 fragments are found to have no more than ten bioisosteres, respectively. Fig. 4 illustrates such an example of modifying an uncommon fragment. The molecular fragment in Fig. 4(a), a 4-azaindole moiety, appears dozens of times in the ChEMBL database as a linker in the visualized fashion. Utilizing our MMP database, only four fragments were found as bioisosteres, and we visualized them in Fig. 4(b). All the found bioisosteres are fused aromatic heterocycles, the same as the original fragment except for one heavy atom, which can form similar non-covalent interactions with biological targets. However, by solely depending on these fragments identified from the MMP analysis, one cannot explore diverse chemical moieties to find potential bioisosteres. More statistics about the number of bioisosteres for each fragment in the fragment library can be found in supplementary figure 3.
@@
To find more bioisosteres beyond them, we employed our model on a molecule from the ChEMBL database by specifying the azaindole moiety as a removal fragment, as in Fig. 4(c). We allowed all the insertion fragments in our fragment library regardless of whether they were used in the training process. Here, to explore bioisosteres of the azaindole moiety that exhibits similar drug-likeness, we set the property control objective as QED=0. With this setting, we generated 50 unique molecular structures from the original molecule with DeepBioisostere.
@@
Fig. 4(d) illustrates eight highly ranked insertion fragments proposed by our model. We noticed that they fall into mainly two types: fused aromatic rings (the orange box) and covalently bonded aromatic rings (the green box). The fragments in the orange box consist of fused aromatic heterocycles, each containing one or two nitrogen atoms, similar to those identified from our MMP database shown in Fig. 4(b). On the other hand, the fragments in the green box have substantially different molecular structures, where each of them contains covalently bonded aromatic rings. Still, all the highly ranked fragments are inserted into the test molecule to exhibit QED values similar to that of the original molecule. (see Fig. 4(d)) Indeed, they contain a few hydrogen bond donors and acceptors and two aromatic rings, the same as the original fragment, and they have a comparable level of lipophilicity. These structural features allow us to expect that they possess the potential to form similar non-covalent interactions as the original one. Fig. 4(e) illustrates the QED distribution of 50 unique modified structures. The peak near the original QED value suggests that most of the 50 insertion fragments are similar to the azaindole moiety in terms of diverse structural features crucial for bio-activity such as the number of hydrogen bond donors and acceptors, aromatic rings, and polar surface area. This example study demonstrates that one can adopt DeepBioisostere for an uncommon chemical moiety to explore potential bioisosteres with a similar level of drug-likeness or QED.
2.5 Targeting SARS-CoV-2 Main Protease Mutant
Lastly, we demonstrate DeepBioisostere in a more practical example, where modifying a known drug to inhibit mutants of the original biological target. The ongoing coronavirus 2019 (COVID-19) pandemic still threatens public health and safety. Due to the limited therapeutic options, a few oral COVID-19 therapeutics have been recently developed, especially for non-hospitalized patients.[43] Ensitrelvir is the first oral noncovalent and non-peptidic SARS-CoV-2 3CL protease () inhibitor that received emergency regulatory approval from the Ministry of Health, Labour and Welfare in Japan.[44, 45] Ensitrelvir shows remarkable biochemical activity and antiviral activity against several mainstream SARS-CoV-2 variants,[44] but recent studies have reported that several naturally occurring mutations of show drug resistance to Ensitrelvir and other oral antiviral drugs.[46, 47, 48] Since these mutations still hold the possibility to undergo changes that show more increased drug resistance, the development of specific therapeutics for them is potentially needed.
We selected the E166V mutant as a target here since it is associated with about 10-fold reduced Ensitrelvir inhibitory activity relative to the wild type. [49] We retrieved and exhibited the complex structures of the wild type(PDB ID: 8HBK) and the E166V mutant(PDB ID: 8H3G) with Ensitrelvir in Fig. 5(a), reported by Duan et al. [49] We analyzed their structures and noticed that they were not much different in their binding pockets. Based on this observation, we hypothesized that a more potent inhibitor on the mutant could be derived from a modification of Ensitrelvir rather than discovering a new lead series.
In the binding structure of the wild type and Ensitrelvir, the E166 residue is known to form multiple H-bondings with another region of protein, contributing to the stability of the overall binding structure. Thus, the change of the E166 residue to Valine must increase the hydrophobicity nearby and interrupt the H-bonding network, resulting in the decreased binding stability of Ensitrelvir or increased drug resistance. [49] In this regard, we hypothesized that modifying the methyl-triazole moiety, circled in gray at Fig. 5(b), of Ensitrelvir to a hydrophobic one would increase sensitivity to the E166V mutant. A large change in molecular size must be avoided during modification since the pocket volume of the E166V mutant is not adequate to contain bulkier fragments. (see Fig. 5(b)) Since Ensitrelvir has been synthesized and marketed as an emergency drug, we expect the synthetic accessibility and drug-likeness to be maintained during bioisosteric replacement. Hence, we set a property control condition as increasing log by 1 while sparing molecular weight, QED, and SAscore to identify an appropriate bioisostere of the methyl-triazole moiety with increased hydrophobicity and maintained size.
We designated the methyl-triazole as a removal fragment to be modified. Adopting the entire fragment library, we obtained 500 modified molecular structures with DeepBioisostere. We scored their binding affinities to the E166V mutant with PIGNet2, a physics-informed deep learning model for predicting protein-ligand binding affinities that exhibits promising performance in comparing binding affinities of structural analogs to the same biological target.[52] We generated conformers of the generated molecules by fixing the conformations of the remaining part, followed by local structural optimization with SMINA,[53] a docking program based on AutoDock Vina.[54] (See supplementary information for details about the overall experimental process.)
The PIGNet2 score distribution of the 500 generated molecules is shown in Fig. 5(c). The average and median scores of all generated molecules were -7.80 kcal/mol and -7.68 kcal/mol, respectively, which are better than the score of the reference ligand, -7.05 kcal/mol. Also, 25.8% of the generated molecules were expected to show 10 times better binding affinity against the E166V mutant than Ensitrelvir in terms of half maximal inhibitory concentration, regarding the difference of 1.36 kcal/mol in energy scale as the 10-fold difference in the inhibitory concentration. We visualized 3D interaction patterns of Ensitrelvir in the binding structure in Fig. 5(d) and that of one generated molecule with a highly ranked affinity score in Fig. 5(e). Notably, the alternative cyclohexanol fragment possesses a broader hydrophobic surface so as to establish extensive hydrophobic interactions with the Val166 residue, while its terminal hydroxyl group also forms an additional hydrogen bond with the Ser1 in another chain. This resulted in a better affinity score against the E166V mutant compared to Ensitrelvir, as intended. Similarly, many molecules with binding affinities better than -9 kcal/mol exhibit extensive hydrophobic interactions with Val166 in addition to an extra hydrogen bond with the Ser1 or Ser144.(see Supplementary Figure 6) Remarkably, along with the enhanced binding affinity scores, the four target properties of the generated molecules consistently satisfy our property control objectives as displayed in Fig. 5(f). This highlights the potential of DeepBioisostere on lead transformation, utilizing bioisosteric replacements to discover better binders while preserving crucial properties in drug development.

### v2

@@ 2.3 Detailed Analysis on Modification with DeepBioisostere
We randomly selected three molecules from the ChEMBL database with moderate molecular weight and lipophilicity, as visualized in Fig. 2(a). We suppose scenario 1 introduced in Section 2.2, where log is to be modulated while a molecular weight is kept unchanged. We generated 100 unique modified molecular structures for each test molecule and property condition with the DeepBioisostere model. Again, only the fragments from the test dataset for insertion were adopted, the same as in the previous section. We exhibited their log values and molecular weights in Fig. 2(b). For every test molecule, the generated structures under different property conditions showed distributions that were completely distinct in terms of log values but very similar in molecular weights. log was increased or decreased by about 1 while most changes in molecular weight were less than 20, corresponding to the weight of one or two heavy atoms.
@@ 2.3.1 Removal Fragment Selection
We looked into the generated molecules and analyzed the removal fragment selection step. We counted how many times each substructure was selected as a removal fragment and visualized them in Fig. 2(c). For the molecule in the middle column, CHEMBL5434005, the carboxylic acid group was most frequently selected as a removal fragment to increase log. In contrast, to decrease log, it was hardly substituted, but the chlorobenzene moiety was preferred as a removal fragment. These choices of removal fragments align well with the known chemical knowledge; introducing a carboxylic acid group enhances hydrophilicity, whereas incorporating an aryl halide moiety increases hydrophobicity. The other two test molecules also showed similar differences in the removal fragment selection under different property control conditions, as exhibited in the 1st and 3rd columns in Fig. 2.
We further visualized three analogs for CHEMBL5434005 generated by DeepBioisostere with the highest likelihoods in Fig. 2(d). When conditioned to increase log (orange arrow), the carboxylic acid group of the parent molecule was replaced with an isopropanol or fluoroethyl group, and the hydroxypyridine scaffold was substituted by a fluorobenzene ring, which are the cases where bioisosteric replacements that introduce more polarity were applied. As a result, the proposed molecular structures exhibited log values higher than the original. Meanwhile, to reduce the log value (blue arrow), our model selected the 4-chlorophenyl moiety on the terminal site of CHEMBL5434005 as a key site for modification. The chosen alternative fragments introduce more hydrophilic aryl groups, reducing the modified structures’ log values. In the last example, the DeepBioisostere model selected a carboxylic acid group, despite its well-known hydrophilicity, as the removal fragment. We believe this is due to the fact that a carboxylic acid is one of the most frequently considered bioisosteres in the hit-to-lead process, which should be reflected in the training data we constructed.[lassalas2016structure] Importantly, our model still succeeded in lowering the log of the given molecule by introducing a more hydrophilic moiety as an alternative. These results suggest that our model’s sequential-decision framework discussed in Section 2.1 effectively balances property-control objectives by choosing appropriate removal fragments and selecting insertion fragments conditioned on diverse combinations.
@@ 2.3.2 Insertion Fragment Selection - Effect of Chemical Environment on Property Control
Another essential factor that governs the selection of the insertion fragments is the chemical environment. Here, we refer to the chemical environment as the substituents surrounding the selected removal substructure to be replaced by a new insertion fragment. We demonstrate its importance in molecular optimization by two single-property control tasks: increasing log by 1 and increasing QED by 0.1.
@@
Fig. 3 illustrates the overall experimental procedure and its results. First, we selected two molecules that contain a pyrazole group, which is a frequently considered aromatic group in bioisosteric replacement.[phenyl-bioisosteres] Their molecular structures are shown in Fig. 3(a); we refer to them as molecules A and B (or and , respectively). The two molecules exhibit great differences in log(5.54 and -0.35) while showing moderate levels of QED values(0.42 and 0.41). Thus, the pyrazole moieties in the two molecules are situated in fundamentally different chemical environments, particularly in terms of the number of aromatic rings, hydrogen bond donors and acceptors, and rotatable bonds. Here, we forced our model to choose only the pyrazole groups as removal fragments to modulate a given target property, resulting in 100 unique modified structures. ( and in Fig. 3(a)) For comparison, we retrieved the insertion fragments selected for molecule A and incorporated them into molecule B, resulting in another set of 100 unique molecules. ( in Fig. 3(a)) In the first two cases, and , the individual chemical environments were directly engaged in the selection of insertion fragments. However, in the case of , the insertion fragments were selected based solely on the chemical environment of molecule A, without any involvement from molecule B. By comparing these cases, we can examine the significance of the chemical environment in molecular property control, especially during the selection of insertion fragments.
@@
Fig. 3(b) and (c) show the property distributions of the modified molecules, where both and nearly achieve the objective values. However, in the case of , the observed property changes in log and QED conditioning tasks were distinct. In the former task, insertion fragments selected to increase log of molecule A also increased log of molecule B, despite the differences in the chemical environments in which the pyrazole moieties were involved. Conversely, in the QED conditioning task, the case of resulted in molecular structures with decreased QED values, while those from and exhibited increased QED values compared to the original ones. This apparently shows that the bioisosteric replacements selected to improve the QED value of molecule A failed to improve that of molecule B. We attribute this distinction to the intrinsic difference between the two properties; log can be readily approximated by summing over the estimated contributions of each chemical moiety in a molecule, while QED is a non-linear function of the molecule-level structural features, such as the numbers of hydrogen bond donors and acceptors. This example highlights the importance of considering the entire chemical environment, not just the removed fragment, when designing a suitable bioisosteric replacement for properties that depend on global molecular context. Such context-dependent property modulation cannot be handled by database-mining approaches, which rely solely on pre-defined fragment statistics. In contrast, DeepBioisostere explicitly learns to capture the non-linear relationships between fragments and their molecular environments, enabling context-aware bioisosteric replacements.
2.4 Exploring Bioisosteres for Uncommon Molecular Fragments
@@
Next, we examine our model in identifying potential bioisosterism. Identification of novel bioisosteres becomes necessary when modification of an uncommon molecular fragment is required. In this case, a typical database-mining approach is likely to be insufficient since only a few bioisosteres from the database might be available for that fragment. Indeed, among 140,193 fragments in the fragment library we created, 98,459 fragments are found to have no more than ten bioisosteres, respectively. Fig. 4 illustrates such an example of modifying an uncommon fragment. The molecular fragment in Fig. 4(a), a 4-azaindole moiety, appears dozens of times in the ChEMBL database as a linker in the visualized fashion. Utilizing our MMP database, only four fragments were found as bioisosteres, and we visualized them in Fig. 4(b). All the found bioisosteres are fused aromatic heterocycles, the same as the original fragment except for one heavy atom, which can form similar non-covalent interactions with biological targets. However, by solely depending on these fragments identified from the MMPA, one cannot explore diverse structural analogues of an acclaimed molecule. More statistics about the number of bioisosteres for each fragment in the fragment library can be found in Supplementary Figure 3.
@@
To find more bioisosteres beyond them, we employed our model on a molecule from the ChEMBL database by specifying the azaindole moiety as a removal fragment, as in Fig. 4(c). We allowed all the insertion fragments in our fragment library regardless of whether they were used in the training process. Here, to explore bioisosteres of the azaindole moiety that exhibit similar drug-likeness, we set the property control objective as QED=0. With this setting, we generated 50 unique molecular structures from the original molecule with DeepBioisostere.
@@
Fig. 4(d) illustrates eight insertion fragments predicted by our model. Under the constraint of QED=0, several fused aromatic rings containing a few hetero-atoms were chosen as bioisosteres of the azaindole moiety. These fragments feature either a six-membered ring fused to a five-membered ring (the green box) or two six-membered rings, in contrast to the original azaindole moiety (the orange box). Despite this variation, all highly ranked fragments, when inserted into the test molecule, exhibit QED values comparable to the original. (see Fig. 4(d)) Each fragment retains a similar number of hydrogen bonding sites, two aromatic rings, and comparable lipophilicity, suggesting that they can engage in analogous non-covalent interactions. Fig. 4(e) illustrates the QED distribution of 50 unique modified structures. The peak near the original QED value implies that most insertions retain key biochemical features, including hydrogen-bond capacity, aromaticity, and polar surface area. This case study demonstrates that DeepBioisostere can be effectively applied to identify potential bioisosteres of uncommon chemical moieties that exhibit drug-likeness profiles comparable to those of the parent moiety.
2.5 In Silico Hit-to-lead Optimization of Structure-based Generative Model-derived Hit Candidates
We further extend the scope of our model to an in silico hit-to-lead workflow, where model-generated hit candidates are refined to improve drug-likeness and synthetic accessibility. Recent 3D molecular generative models for structure-based drug design have focused more on binding affinity to a given target protein, especially considering the structural context of the pocket environment in which the interaction between a protein and a generated molecule occurs.[pocket2mol, deepicl, targetdiff, decompdiff] Despite the target-specific design capacity of these approaches, generated molecules often fail to meet the criteria of experimental hits due to other molecular properties.
This gap between computational and actual hits highlights the need for a robust strategy to refine model-generated hit candidates, or in other words, to enhance molecular properties while preserving binding affinities. Hence, we demonstrate the synergy of using DeepBioisostere for multi-property optimization on top of target-specifically designed molecules from molecular generative models, revealing the applicability of our strategy. The goal here is to optimize synthesizability and drug-likeness of hit candidates, estimated as SAscore and QED, respectively, without losing their bioactivities towards respective targets.
We initially obtained computational hit candidates by four recent state-of-the-art structure-based 3D molecular generative models: Pocket2Mol[pocket2mol], DeepICL[deepicl], TargetDiff[targetdiff], and DecompDiff[decompdiff]. From the CrossDocked2020 test set, we selected three target proteins whose generated molecules exhibit both poor drug-likeness and synthetic accessibility, making them suitable targets that can take the most advantage from chemical modification. The details of the model used and the criteria for target selection, along with the statistics of the generated molecules, are provided in the Supplementary Information Section 7. Generated molecules for each target and model, termed computational hit candidates, underwent modifications through different replacement selection strategies, including DeepBioisostere. For comparison of our model with conventional database-mining methods, we adopted three baseline replacement strategies, which are referred to as random, frequency-based, and MMPA-based, as shown in Table 2. The three strategies were set to utilize the bioisosteric pairs in the MMP database we constructed. In particular, the MMPA-based strategy first filters out replacements whose occurrence in our database is less than 10 times, and then selects replacements that exhibited the highest improvement in both QED and SAscore in the database. This process is based on statistical evidence in the MMP database, in the same way with the conventional database-mining strategies, such as SwissBioisostere. Here, we guide DeepBioisostere to increase QED by 0.1 and to decrease SAscore by 1.0, simultaneously. For each method, replacement was attempted 100 times for each input molecule.
| Model | Strategy | Success Rate | |||
| QED | SAscore | Docking | Joint | ||
| DeepICL | Random | 0.19 | 0.17 | 0.97 | 0.06 |
| Frequency-based | 0.30 | 0.26 | 0.99 | 0.10 | |
| MMPA-based | 0.33 | 0.22 | 0.97 | 0.11 | |
| DeepBioisostere | 0.69 | 0.80 | 0.96 | 0.56 | |
| Pocket2Mol | Random | 0.62 | 0.39 | 0.91 | 0.24 |
| Frequency-based | 0.67 | 0.51 | 0.97 | 0.36 | |
| MMPA-based | 0.64 | 0.42 | 0.98 | 0.30 | |
| DeepBioisostere | 0.75 | 0.74 | 0.94 | 0.56 | |
| TargetDiff | Random | 0.32 | 0.50 | 0.96 | 0.18 |
| Frequency-based | 0.43 | 0.58 | 0.98 | 0.25 | |
| MMPA-based | 0.42 | 0.50 | 0.99 | 0.20 | |
| DeepBioisostere | 0.83 | 0.87 | 0.97 | 0.74 | |
| DecompDiff | Random | 0.38 | 0.39 | 0.95 | 0.19 |
| Frequency-based | 0.49 | 0.45 | 0.98 | 0.24 | |
| MMPA-based | 0.43 | 0.29 | 0.99 | 0.13 | |
| DeepBioisostere | 0.79 | 0.87 | 0.96 | 0.68 | |
For evaluation, we compared the success rates of QED and SAscore optimization, as well as the retention of docking scores to a protein target. The success criteria for QED and SAscore are based on whether the generated molecule exhibits better properties than the original. We consider a generated docking score as a success when it is retained during replacement within 1.36 kcal/mol, corresponding to a 10-fold change in IC50. The most integral metric, which is termed ‘Joint’, means the ratio of generated molecules jointly meets all the aforementioned criteria—improved QED, SAscore, and preserved docking score—thereby indicating the success in the computational hit-to-lead scenario for increasing synthesizability and drug-likeness.
Table 2 shows the result of a target protein with PDB ID 4RV4, whereas the other results are shown in the Supplementary Table 4. Although frequency- and MMPA-based strategies performed well at docking score retention—possibly as a result of the use of our bioisosteric replacement library curated from bio-assays—DeepBioisostere mostly achieved superior performance in optimizing the properties such as QED and SAscore. We attribute the improved controllability of QED and SAscore to the consideration of the entire molecular graph, along with the removal and insertion fragments, rather than focusing solely on a single fragment as in the baselines. The true power of DeepBioisostere is demonstrated in its ability to satisfy all criteria simultaneously, as evidenced by its superior performance in the ‘Joint’ metric, for which baseline approaches exhibited less than 10 % success.
Figure 5 shows two example molecules generated based on bioisosteric replacements identified by DeepBioisostere, both of which achieve improved QED and SAscore while preserving the binding affinity of the original molecule. Notably, these improvements arise from the substitution of the fragments resulting in amide-containing molecule , which is expected to enhance the synthetic accessibility. Together, this robust ability to find replacements that fulfill multiple objectives is notably agnostic to underlying generative models, establishing it as a broadly applicable and powerful strategy for in silico drug design.